OpenStudy (anonymous):

The perimeter of a carpet is 90 feet. The width is 11 feet more than the length. Find the width of the carpet. 17 feet 18 feet 24 feet 30 feet

OpenStudy (anonymous):

first you need to go down your answer choices until you get the right one

rebeccaxhawaii (rebeccaxhawaii):

or instead of wasting time

rebeccaxhawaii (rebeccaxhawaii):

do you have the formula

OpenStudy (anonymous):

No I do not have the formula that is all I was provided when given the question

OpenStudy (anonymous):

well the formula is x-11=y

OpenStudy (anonymous):

y+blank+y+blany=perimiter

OpenStudy (anonymous):

so for example I would do y+17+y+17=90 and i keep going through all the choices until one is right?

OpenStudy (sagewilson):

Yeah pretty much.

rebeccaxhawaii (rebeccaxhawaii):

still confused?

OpenStudy (anonymous):

yes clearly im not college math material

OpenStudy (anonymous):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

17+17+28+28=90

OpenStudy (anonymous):

ohhhhh so i use this formula every time i have a question similar to this one